Fermi National Laboratory


Accelerator Update

Friday October 24 through Monday October 27
All shifts began in shutdown.

On Friday, through out the day electricians continued their Tevatron (TeV) service building electrical panel maintenance work. During the morning, high voltage experts took down feeder 46-A for about an hour for maintenance. In the afternoon, operators interlocked MI-13 in the Main Injector (MI) enclosure. During the evening shift, operators interlocked TeV sector F and the Booster and also successfully conducted a highpot.

On Saturday, during the midnight shift, operators made Booster and MI ready for beam by exercising as many systems as possible. Of course many things tripped off and were reset, but a Booster Klystron Debuncher tripped off due to a high temperature and Operations decided to leave it off for further investigation in the morning. Operators put MI into access mode near the end of the midnight shift; during the highpot portion of this procedure operators discovered a ground fault. High voltage experts took down feeder 45 for maintenance; this was an eight-hour job. The day shift operators made the Booster ready for access. Late in the afternoon a MI expert reported finding a bad quadrupole magnet (Q206). During the evening shift operators interlocked the Antiproton Source (Pbar).

On Sunday, the midnight shift operators exercised Booster devices. Near the end of the shift a RF station (BRF13) went down hard. Operators contacted RF technicians. Operators also found a problem with a Booster vacuum valve. During the day shift an expert turned on Pbar power supplies. He discovered a ground fault; he and operators accessed and cleared the fault. Linac developed data logger problems around noon. Safety system techs tested systems throughout the evening shift.

On Monday, midnight shift operators prepared areas for normal access and work.

WEEKEND STATUS SUMMARY
Linac

Booster
  • Vacuum work at sector long-4 was completed and now pumping down
  • BRF13 has a bias/screen fault
  • B:MP02X is up and running
  • Booster Kickers are on
  • B:BEX3 has a klixon problem
  • B:Q802 has a safety system problem
  • B:MI18BD is up an running
  • DOGL3 won't turn on
  • B:MV1 and B:VBC1 locked off by RSO
  • B:MH10 is off and B:MV10 is on
  • CES7, CEL7, and CES17 all have readback problems
  • All power supplies not mentioned above are running

MiniBooNE

  • CDC input cleared
  • E:HV860 locked off by RSO
  • All enclosures secure
Main Injector
  • Q206 has a ground fault
  • All other ground fault detection checked okay
  • I:KPS1C filament is not ready (It will be looked at on 10/27/03)
  • Transfer lines P1 and A1 have had the power supplies turned on
    • I:HVF12 and I:Q703 need work
    • I:HVF11 may have a ground fault in tunnel
    • S:HP3DS is causing problems with the P1 CDC

Tevatron

  • LCW systems are on and all pond pumps returned to pre-power outage state
    • Sector E1 LCW readbacks are bad
  • All Vacuum systems are on
  • The CAMAC crate files have been restored

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/27/2003   email Fermilab